# From Turbulence to Tranquility: AI-Driven Low-Altitude Network

### v1

@@ Abstract
Low Altitude Economy (LAE) networks own transformative potential in urban mobility, emergency response, and aerial logistics. However, these networks face significant challenges in spectrum management, interference mitigation, and real-time coordination across dynamic and resource-constrained environments. After addressing these challenges, this study explores three core elements for enabling intelligent LAE networks as follows machine learning–based spectrum sensing and coexistence, artificial intelligence (AI)-optimized resource allocation and trajectory planning, and testbed-driven validation and standardization. We highlight how federated and reinforcement learning techniques support decentralized, adaptive decision-making under mobility and energy constraints. In addition, we discuss the role of real-world platforms such as AERPAW in bridging the gap between simulation and deployment and enabling iterative system refinement under realistic conditions. This study aims to provide a forward-looking roadmap toward developing efficient and interoperable AI-driven LAE ecosystems.
@@
Low-altitude economy, unmanned areal vehicle (UAV) communications, machine learning, resource optimization.
@@ I Introduction
Low Altitude Economy (LAE) network has emerged as one of the most technological trends of the next decade. Defined as using airspace below $30003000$ meters for economic, social, and operational activities, LAE covers various applications: urban air mobility such as air taxis, emergency medical deliveries, precision agriculture, environmental sensing, surveillance, and logistics. Supported by unmanned aerial vehicles (UAVs), electric vertical takeoff and landing (eVTOL) aircrafts, high-altitude balloons, and drone swarms, LAE holds enormous promise to revolutionize transportation systems, reduce urban congestion, improve service delivery, and create new economic opportunities. LAE’s importance can be extended beyond pure innovation by addressing critical global challenges such as climate change mitigation, urban sustainability, and disaster resilience [1]. Therefore, LAE can offer an unprecedented opportunity to transform how cities operate and how services are delivered in both urban and remote areas since it enables faster, cleaner, and more flexible aerial operations.
@@
The integration of these networks with terrestrial systems adds further complexity. To realize the full potential of LAE, seamless coexistence between airborne, terrestrial, and non-terrestrial networks should be provided. Thus, this requires new strategies for dynamic spectrum access, adaptive channel modeling, robust interference management, and collaborative multi-layer network control [2]. Due to complex system architecture, needs, and stringent requirements, traditional rule-based approaches can pose indeadequate to address these challenges. Instead, deep learning (DL) based approaches can play a significant role in meeting the requirements for safe, efficient, and adaptive LAE networks.
@@
DL, particularly distributed learning, offers powerful tools to enable harmonious coexistence of air and ground networks. With federated learning, UAVs can collaboratively sense spectrum usage without central coordination [3]. Reinforcement learning enables aircraft to dynamically adjust their trajectories and resource allocations in response to environmental changes. Generative models such as conditional generative adversarial networks (CGANs) help predict radio conditions and optimize channel utilization [4]. Orchestration of these DL techniques promises to make LAE networks more adaptable, resilient, and scalable.
We introduce a comprehensive architecture designed to integrate communication, computing, and sensing functions in LAE networks, which envisions an intelligent control layer that coordinates air and ground assets using AI-powered models as illustrated in Fig. 1. Combining dynamic spectrum management, edge computing, real-time spectrum sensing, and multi-agent coordination, it aims to create an ecosystem where various platforms can share airspace safely and efficiently. To be able to run such an ecosystem, coordination across layers is essential, particularly in environments experiencing spectrum congestion and high mobility.
In this study, we first investigate challenges in terms of wireless communications and sensing towards fully operational LAE networks and then three fundamental elements for advancing LAE: (i) ML for spectrum sensing and coexistence (ii) ML-assisted resource optimization and trajectory planning, and (iii) the development of standards and real-world validation using our testbeds. For each pillar, we provide a detailed review of open issues, current progress, possible solutions using DL, and future directions. Our ultimate goal is to present a roadmap that bridges the gap between theoretical innovation and practical deployment, enabling the realization of scalable, secure, and intelligent LAE networks.
@@ II Challenges in Low-Altitude Wireless Networks
The growth in LAE systems introduces significant challenges across multiple technical aspects. Although numerous challenges exist, such as air traffic management, energy limitations, security risks, privacy concerns, and regulatory compliance [2], this study focuses specifically on spectrum coexistence, integration of communications, computing, and sensing (C2S), and standardization and testbed, which are addressed in subsections below.
@@ II-A ML in Spectrum Utilization and Coexistence
Due to the highly dynamic and dense nature of low-altitude airspace, optimization of spectrum utilization poses formidable challenges. For example, high mobility among UAVs and eVTOLs leads to Doppler shifts, frequent handoffs, and time-varying multipath fading, which significantly impacts link stability. Another issue is that interference becomes particularly inevitable in urban environments where aerials must coexist with terrestrial base stations, Wi-Fi networks, legacy aviation systems, and radars. Therefore, how to dynamically sense and allocate spectrum resources, predict interference patterns in real time, and balance among competing air and ground users are the main questions remaining open, yet [5].
@@
On the other hand, ML offers opportunities to optimize spectrum utilization and interference management in LAE networks; however, its practical application requires solid solutions for several challenges. One of the most important issues is the non-stationary and spatially correlated nature of the environment. Unlike terrestrial networks, LAE consists of highly mobile platforms such as UAVs and eVTOLs operating at different altitudes and mediums. As a result, it leads to changing signal propagation conditions, including frequent transitions between line-of-sight (LOS) and non-line-of-sight (NLOS) links, and fluctuations in interference levels. ML models trained on localized or static datasets have a risk of failing to generalize across the entire 3D airspace. This might reduce the reliability of spectrum estimation and classification.
@@
Another important challenge is the scarcity of labeled and reliable data required to train robust spectrum sensing models. In LAE environments, acquiring ground truth information about spectrum occupancy or interference is not only costly, but also sometimes impractical. For example, military radar systems or special emergency broadcasts may be difficult to detect or label accurately.
@@
Moreover, cross-layer dynamics create additional complexity. Spectrum utilization in LAE is inherently a cross-layer problem since physical layer measurements such as signal-to-noise ratio (SNR) and Doppler shift interact with medium access control (MAC) and network layers. ML models trained in isolation without timely feedback from other layers may become out of sync with the operational context, which is known as model drift [6]. In practice, synchronization between layers is difficult to achieve because of limited cross-platform communication and the asynchronous nature of distributed LAE nodes. This can reduce spectrum reuse efficiency or harm coordination in dynamic spectrum access.
@@
Finally, the lack of standardized simulation environments and datasets tailored to LAE-specific conditions significantly hinders progress. Most existing ML models for spectrum estimation or interference mitigation are evaluated under simplified or terrestrial scenarios that fail to capture the 3D, multi-agent, and time-varying characteristics of low-altitude networks. In the absence of common metrics, reproducibility is limited, and the generalizability of findings remains uncertain. Therefore, regulation and standardization efforts are essential to establish a reliable ML-based spectrum management infrastructure. Moreover, the opaque decision-making processes of many ML models limit their reliability and regulatory acceptance, particularly in spectrum-sensitive domains such as aviation, where explainability and robustness are prerequisites for deployment. As a result, the integration of ML into LAE spectrum systems is likely to remain an open challenge until these domain-specific constraints are effectively addressed.
@@ II-B Integration of C2S
A critical challenge is the scheduling of C2S tasks across a swarm of heterogeneous aerial nodes. For instance, high-resolution sensing workloads can saturate data pipelines so that communication bottlenecks can delay time-sensitive decisions such as path changes or mission replanning. The lack of unified control systems might cause suboptimal performance and poor resource utilization. Furthermore, heterogeneity in UAV platforms (e.g., varying in computational capacity, sensor types, and antenna systems, etc) makes synchronized data fusion and task assignment further challenging [2].
Finally, a key challenge is the lack of standardized architectures that can support scalable, fault-tolerant C2S integration. Current implementations tend to stack functions vertically rather than holistically designing coordinated workflows with feedback across layers. Bridging this gap requires not only algorithmic developments but also robust system designs that consider the interactions among the physical layer, the network, and sensors.
@@ II-C Standardization and Testbed
In parallel, the limited availability of publicly available testbeds undermines the validation and benchmarking of proposed systems and algorithms. Although simulation remains useful for preliminary analysis, it fails to capture some physical layer nuances of LAE environments, such as altitude-dependent multipath fading, spatiotemporal spectrum utilization, and platform-dependent distortions. As detailed later, to address this, the Aerial Experiments and Research Platform for Advanced Wireless (AERPAW) testbed has emerged as a critical resource enabling controlled experiments with programmable UAVs and software-defined radios (SDRs) in mixed urban-rural landscapes [8]. AERPAW supports the evaluation of AI-driven spectrum access and cross-layer optimization under realistic constraints, bridging theory and practice, and hosts over two dozen datasets based on real-world UAV network environments.
@@ III AI-Driven Spectrum Management and Interference Mitigation
Figure 2: Federated learning based spectrum sensing performance results versus (a) transmit power, (b) number of UAVs, and (c) training data per UAV.
Spectrum management and interference detection constitute key enablers for robust LAE network operations. Since aerial systems become denser and more dynamic, the ability to intelligently sense, allocate, and manage spectrum resources is essential. AI techniques, particularly ML, have emerged as powerful tools to address these challenges, offering scalable, adaptive, and data-driven approaches. Accordingly, this section discusses ML-based approaches for spectrum sensing and interference detection.
@@ III-A AI-Enabled Dynamic Spectrum Sensing for Low-Altitude Networks
Dynamic spectrum sensing is a fundamental component for enabling reliable and efficient communication in the highly congested LAE, where aerial and terrestrial users share limited frequency resources. In [3], we proposed an federated learning (FL) method to address the challenges of distributed UAV networks, where centralized solutions are impractical because of power constraints, privacy concerns, and high communication overhead. The proposed system utilizes a lightweight convolutional neural network (CNN) architecture, which is specifically designed to operate on resource-constrained UAV platforms. By processing the real and imaginary components of received signals using one-dimensional convolutions, the system captures temporal signal patterns while trying to maintain computational efficiency. A key contribution of the study is the introduction of the FedSNR aggregation method, extending the conventional federated averaging (FedAvg) approach by incorporating the SNR observed at each UAV into the aggregation process. This approach allows the global model to give greater weight to updates from UAVs experiencing higher-quality signals, thereby enhancing adaptability, which is particularly important in heterogeneous and dynamic environments.
[…]
@@
In summary, the proposed FL framework for UAV-based spectrum sensing offers a privacy-preserving and communication-efficient solution. Also, it significantly improves detection accuracy. By joint utilization of a distributed lightweight CNN model with FedSNR aggregation strategy, the system addresses accurate sensing and resource constraints in LAE networks. These findings provide an important foundation for future research and practical implementations in real-world aerial communication systems, where spectrum-sharing requirements demand adaptive and efficient learning solutions.

### v2

# AI-Driven Low-Altitude Economy: Spectrum, Mobility, and Validation

@@ Abstract
The Low Altitude Economy (LAE) network, with its transformative capabilities, is a candidate to become one of the major technological developments of the next decade for air mobility. However, the expected unprecedented density, mobility, and heterogeneity pose challenges and require new approaches, as it renders traditional rule-based approaches inadequate. To address these challenges, this study introduces artificial intelligence (AI)-based approaches and validation frameworks for transitioning AI-enabled technologies from simulation-based studies to practical and deployable systems. This study discusses essential enablers for intelligent LAE networks. First, AI-based spectrum sensing and coexistence utilizing the distributed nature of LAE nodes is introduced. Then, joint resource allocation and trajectory optimization driven by reinforcement learning is discussed. Bridging the gap between simulation and deployment through experimental platforms such as Aerial Experiments and Research Platform for Advanced Wireless (AERPAW), which are critical for validating models under realistic and non-stationary airspace conditions, is also addressed. The study concludes by highlighting open issues and outlining a forward-looking roadmap for the development of efficient, interoperable, and scalable AI-driven LAE ecosystems.
@@
Low-altitude economy, unmanned aerial vehicle (UAV) communications, machine learning, resource optimization.
@@ I Introduction
The Low Altitude Economy (LAE) network is poised to become one of the defining technological trends of the next decade. Encompassing the use of the airspace below $3000$ metres for economic, social, and operational activities, LAE covers various applications: urban air mobility (e.g., air taxis, emergency medical deliveries), precision agriculture, environmental sensing, surveillance, and logistics, as illustrated ixn Fig. 1. Supported by unmanned aerial vehicles (UAVs), electric vertical takeoff and landing (eVTOL) aircraft, high-altitude balloons, and drone swarms, LAE holds enormous promise to revolutionize transportation systems, reduce urban congestion, improve service delivery, and create new economic opportunities. LAE’s importance can be extended beyond pure innovation by addressing critical global challenges such as climate change mitigation, urban sustainability, and disaster resilience [4]. Therefore, LAE can offer an unprecedented opportunity to transform how cities operate and how services are delivered in both urban and remote areas because it enables faster, cleaner, and more flexible aerial operations.
I-A Motivations
@@
Integration of these networks with terrestrial systems adds further complexity. To realize the full potential of LAE, seamless coexistence between airborne, terrestrial, and non-terrestrial networks should be provided. Thus, this requires new strategies for dynamic spectrum access, adaptive channel modelling, robust interference management, and collaborative multi-layer network control [5]. Due to the complex system architecture, needs, and stringent requirements, traditional rule-based approaches can be inadequate for addressing these challenges. Instead, deep learning (DL)-based approaches can play a significant role in meeting the requirements for safe, efficient, and adaptive LAE networks. Combining dynamic spectrum management, edge computing, real-time spectrum sensing, and multi-agent coordination, it aims to support an ecosystem where diverse platforms can share airspace safely and efficiently. To operate such an ecosystem, coordination across layers is essential, particularly in environments experiencing spectrum congestion and high mobility.
@@
DL offers powerful tools such as distributed and generative learning strategies to enable the harmonious coexistence of air and ground networks. With federated learning (FL), UAVs can collaboratively sense spectrum usage without central coordination [10]. Reinforcement learning enables aircraft to dynamically adjust their trajectories and resource allocations in response to environmental changes. Generative models such as conditional generative adversarial networks (CGANs) help predict radio conditions and optimize channel utilization [8]. The orchestration of these DL techniques promises to make LAE networks more adaptable, resilient, and scalable.
I-B Research Focus
Despite the promising results that DL techniques have shown in theoretical analysis, simulations, or controlled artificial intelligence (AI) environments using synthetic data, numerous open questions remain before these methods can be effectively implemented in real-world LAE operations. Practical deployments often face challenges such as dynamic spectrum environments, mobility constraints, limited on-board computation, and communication delays. Therefore, this study is dedicated to bridging this gap by examining both the most promising DL-based approaches and the persistent challenges that hinder their practical implementation. In particular, we emphasize the importance of realistic experimentation and validation through advanced testbeds and digital twin platforms such as Aerial Experiments and Research Platform for Advanced Wireless (AERPAW), which represent a critical step towards operationalizing intelligent, resilient, and scalable LAE networks. Fig. 2 shows the framework illustrating key challenges in LAE networks, AI-enabled solutions, and their experimental validation. The challenges towards practical LAE systems are introduced in Section II. Spectrum scarcity and dynamic interference management motivate distributed spectrum sensing addressed in Section III. Also, high mobility and 3D heterogeneous networks require adaptive integration of communications, computing, and sensing (C2S) together with trajectory-aware optimization discussed in Section IV. Finally, due to the lack of realistic datasets and standardized evaluation environments, empirical validation is a must, which is the core of Section V through testbed-driven experimentation and data collection. This challenge–enabler–validation loop proposes a systematic roadmap for transitioning AI-enabled LAE technologies from simulation-based studies to practical and deployable systems.
Figure 2: Framework illustrating the interaction between LAE challenges, AI enablers, and experimental validation toward deployable intelligent aerial networks.
@@ II Challenges in Low-Altitude Wireless Networks
The growth in LAE systems introduces significant challenges across multiple technical aspects. Although numerous challenges exist, such as air traffic management, energy limitations, security risks, privacy concerns, and regulatory compliance [5], this study focuses specifically on spectrum coexistence, integration of C2S, and standardization and testbed, which are addressed in subsections below.
@@ II-A ML in Spectrum Utilization and Coexistence
Due to the highly dynamic and dense nature of low-altitude airspace, optimizing spectrum utilization presents major challenges. For example, high mobility among UAVs and eVTOLs leads to Doppler shifts, frequent handoffs, and time-varying multipath fading, significantly degrading link stability. Another issue is that interference becomes particularly inevitable in urban environments where aerial vehicles must coexist with terrestrial base stations, Wi-Fi networks, legacy aviation systems, and radars. Therefore, how to dynamically sense and allocate spectrum resources, predict interference patterns in real time, and balance among competing air and ground users are the main questions remaining open, yet.
@@
ML offers opportunities to optimize spectrum utilization and interference management in LAE networks; however, its practical application requires solid solutions for several challenges. One of the most important issues is the non-stationary and spatially correlated nature of the environment. Unlike terrestrial networks, LAE consists of highly mobile platforms such as UAVs and eVTOLs operating at different altitudes and media. As a result, it leads to changing signal propagation conditions, including frequent transitions between line-of-sight (LoS) and non-line-of-sight (NLoS) links, and fluctuations in interference levels. ML models trained on localized or static datasets have a risk of failing to generalize across the entire 3D airspace. This might reduce the reliability of spectrum estimation and classification.
@@
Another important challenge is the scarcity of labelled and reliable data required to train robust spectrum sensing models. In LAE environments, acquiring ground truth information about spectrum occupancy or interference is not only costly but also sometimes impractical. For example, military radar systems or special emergency broadcasts may be difficult to detect or label accurately.
@@
Moreover, cross-layer dynamics create additional complexity. Spectrum utilization in LAE is inherently a cross-layer problem because physical layer measurements, such as signal-to-noise ratio (SNR) and Doppler shift, interact with medium access control (MAC) and network layers. ML models trained in isolation without timely feedback from other layers can become out of sync with the operational context, which is known as model drift [6]. In practice, synchronization between layers is difficult to achieve because of limited cross-platform communication and the asynchronous nature of distributed LAE nodes. This can reduce spectrum reuse efficiency or harm coordination in dynamic spectrum access.
@@
Finally, the lack of standardized simulation environments and datasets tailored to LAE-specific conditions significantly delays progress. Most existing ML models for spectrum estimation or interference mitigation are evaluated under simplified or terrestrial scenarios that fail to capture the 3D and time-varying characteristics of low-altitude networks. In the absence of common metrics, reproducibility is limited, and the generalizability of findings remains uncertain. Therefore, regulation and standardization efforts are essential to establish a reliable ML-based spectrum management infrastructure. Moreover, the blackbox decision-making processes of many ML models limit their reliability and regulatory acceptance, particularly in spectrum-sensitive domains such as aviation, where explainability and robustness are prerequisites for deployment. As a result, the integration of ML into LAE spectrum systems is likely to remain an open challenge until these domain-specific constraints are effectively addressed. As we discuss later, digital twin systems that mimic the real environment and experiment platforms are the key enablers for developing and testing these kinds of methods. They accelerate and make the progress toward practical ML model deployments more efficient.
@@ II-B Integration of C2S
A critical challenge is the scheduling of C2S tasks across a swarm of heterogeneous aerial nodes. For instance, high-resolution sensing workloads can saturate data pipelines so that communication bottlenecks can delay time-sensitive decisions such as path changes or mission replanning. The lack of unified control systems might cause suboptimal performance and poor resource utilization. Heterogeneity in UAV platforms, e.g., varying in computational capacity, sensor types, and antenna systems, etc., makes synchronized data fusion and task assignment further challenging [5].
Finally, a key challenge is the lack of standardized architectures that can support scalable, fault-tolerant C2S integration. Current implementations tend to stack functions vertically rather than holistically design coordinated workflows with feedback across layers. Bridging this gap requires not only algorithmic developments but also robust system designs that consider the interactions among the physical layer, the network, and sensors [13]. Along with these, proper test environments accurately simulating real-world scenarios, including the interactions between the physical layer and upper layers as well as hardware characteristics, are required.
@@ II-C Standardization and Testbed
Deployment of LAE networks requires a complex standardization landscape that imposes significant challenges. Therefore, regulators such as Ofcom aim to establish trust and connectivity by shifting generic license exemptions to specific UAV radio licenses, thereby creating interference-protected airspaces [9]. Also, the Federal Aviation Administration (FAA) introduces urban air mobility corridors to manage air traffic with AI agents learning from dynamic and community-based rules rather than rigid clearances. For the full integration and autonomy, the International Civil Aviation Organization (ICAO) proposes a new framework that enables AI systems handle frequent decisions currently limited by human reaction times to harmonize altitude and separation standards globally [12]. However, these visions require experimental validation, yet the physical infrastructure to bridge the gap between these conceptual standards and deployment remains underdeveloped.
@@
In parallel, the limited publicly available testbeds undermine the validation and benchmarking of proposed systems and algorithms. Although simulation remains useful for preliminary analysis, it fails to capture some physical layer nuances of LAE environments, such as altitude-dependent multipath fading, spatiotemporal spectrum utilization, and platform-dependent distortions. As detailed later, to address this, the AERPAW testbed has emerged as a critical resource enabling controlled experiments with programmable UAVs and software-defined radios (SDRs) in mixed urban-rural landscapes [7]. AERPAW supports the evaluation of AI-driven spectrum access and cross-layer optimization under realistic constraints, bridging theory and practice, and hosts over two dozen datasets based on real-world UAV network environments.
@@ III AI-Driven Spectrum Management and Interference Mitigation
(a) Transmit power
(b) Dataset size per UAV
Figure 3: FL-based spectrum sensing performance results.
Spectrum management and interference detection constitute key enablers for robust LAE network operations. Because aerial systems become denser and more dynamic, the ability to intelligently sense, allocate, and manage spectrum resources is essential. AI techniques, particularly ML, have emerged as powerful tools to address these challenges, offering scalable, adaptive, and data-driven approaches. Accordingly, this section discusses ML-based approaches for spectrum sensing and interference detection.
@@ III-A AI-Enabled Dynamic Spectrum Sensing for Low-Altitude Networks
Dynamic spectrum sensing is a fundamental component for enabling reliable and efficient communication in highly congested LAE, where aerial and terrestrial users share limited frequency resources. In [10], we proposed an FL method to address the challenges of distributed UAV networks, where centralized solutions are impractical because of power constraints, privacy concerns, and high communication overhead. The proposed system utilizes a lightweight convolutional neural network (CNN) architecture, which is specifically designed to operate on resource-constrained UAV platforms. By processing the real and imaginary components of received signals using one-dimensional convolutions, the system captures temporal signal patterns while trying to maintain computational efficiency. A key contribution of the study is the introduction of the FedSNR aggregation method, extending the conventional federated averaging (FedAvg) approach by incorporating the SNR observed at each UAV into the aggregation process. This approach allows the global model to give greater weight to updates from UAVs experiencing higher-quality signals, thereby enhancing adaptability, which is particularly important in heterogeneous and dynamic environments.
The experimental setup in [10] involves UAV networks operating in a realistic 3D space, with radar systems generating various waveform types, including continuous wave, chirp, and pulse signals. The CNN is carefully designed with batch normalization and dropout layers to improve training stability and generalization. A global average pooling layer is employed to reduce model complexity and memory requirements, enabling deployment on resource-constrained UAVs on edge devices. Numerical simulations demonstrate the proposed system, with FedSNR achieving up to a $15\%$ improvement in sensing accuracy over centralized models [10]. Notably, as the number of UAVs increased, the system’s sensing accuracy improves. It shows the benefit of incorporating multi-channel spatial diversity in the learning process. Although the proposed approach outperforms FedAvg, particularly in low SNR regimes as indicated in Fig. 3(a), we further test the framework by adding error to the SNR used in FedSNR aggregation because gathering the actual SNR might be challenging in practice, and the estimated SNR might include error, of course. To simulate this kind of error-prone SNR estimation, we feed the SNR values into the aggregation process by adding Gaussian noise, $\mathcal{N}(0,\sigma^{2})$, to the real SNR. As seen in Fig. 3(a), erroneous SNR estimations hinder FedSNR’s learning, especially at low transmit power. Thus, these results point to a gap in the use of the FedSNR method in practice and require a good SNR estimator with an error around $\sigma=1$ dB. Then, Fig. 3(b) shows a critical issue in distributed learning when data scarcity causes a weight divergence problem because environmental conditions and mobility patterns vary widely between nodes, yielding generally a non-independent and identically distributed (non-IID) dataset. This situation highlights why the real environment and digital twin validation are sine qua non for distributed systems, because the simulations might overlook the impact of SNR estimation error, mobility, and non-IID data distributions on the system performance. However, these practical challenges lead to model divergence and degrade performance under data scarcity. As a result, this gap remains open, and addressing it requires the use of experimental platforms like AERPAW, as we discuss in the following sections. These should be carefully investigated in future studies.
@@
In summary, the proposed FL framework for UAV-based spectrum sensing offers a privacy-preserving and communication-efficient solution. By joint utilization of a distributed lightweight CNN model with FedSNR aggregation strategy, the system addresses accurate sensing and resource constraints in LAE networks. However, the issues discussed above indicate crucial open points towards practical applications. These outcomes provide an important insight for future research and practical implementations in real-world aerial communication systems, where spectrum-sharing requirements demand adaptive and efficient learning solutions.